### v1

# Learning Frequency-Aware Dynamic Transformers for All-In-One Image Restoration

Abstract
This work aims to tackle the all-in-one image restoration task, which seeks to handle multiple types of degradation with a single model. The primary challenge is to extract degradation representations from the input degraded images and use them to guide the model’s adaptation to specific degradation types. Recognizing that various degradations affect image content differently across frequency bands, we propose a new all-in-one image restoration approach from a frequency perspective, leveraging advanced vision transformers. Our method consists of two main components: a frequency-aware Degradation prior learning transformer (Dformer) and a degradation-adaptive Restoration transformer (Rformer). The Dformer captures the essential characteristics of various degradations by decomposing inputs into different frequency components. By understanding how degradations affect these frequency components, the Dformer learns robust priors that effectively guide the restoration process. The Rformer then employs a degradation-adaptive self-attention module to selectively focus on the most affected frequency components, guided by the learned degradation representations. Extensive experimental results demonstrate that our approach outperforms the existing methods on four representative restoration tasks, including denoising, deraining, dehazing and deblurring. Additionally, our method offers benefits for handling spatially variant degradations and unseen degradation levels.
Index Terms:
All-in-one Image Restoration, Frequency-Aware Learning, Vision Transformers.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Image restoration aims to reconstruct high-quality images from degraded ones affected by issues like noise, blur, resolution loss, and various corruptions. Over time, this field has found extensive applications in diverse real-world scenarios, spanning general visual perception, medical imaging, and satellite imaging. Prevailing image restoration efforts center on the meticulous design of task-specific approaches and have demonstrated promising results in tasks such as denoising [1, 2, 3, 4], deraining [5, 6, 7, 8], and deblurring [9, 10, 11, 12]. Despite their success in specific tasks, these approaches often prove inadequate when faced with changes in the degradation task or its severity. This limitation presents significant challenges to their practical use in real-world situations, especially in complex environments. For instance, self-driving cars may encounter consecutive or simultaneous challenges, such as rainy and hazy weather. Consequently, it becomes imperative to develop more generalized approaches capable of recovering images from a variety of unknown degradation types and levels.
Recent studies, e.g., [13, 14], have tried to handle multiple degradations with a multitask learning framework. This involves processing images with different types of degradation by sharing a common backbone and designing task-specific heads. Despite the success of multitask methods in image restoration, those with shared parameters often face the challenge of task interference and still require degradation prior during testing. To avoid these drawbacks, all-in-one image restoration has been studied recently, pioneered by Li et al. [15]. This task aims to address various degradation tasks within a single model. Within the all-in-one framework, the crucial problem to be tackled is how to obtain the degradation representations from the degraded images and how to use the acquired degradation representations in the restoration network. In this work, we propose a new approach to tackle the aforementioned challenges by leveraging advanced vision transformers and recognizing that different degradations impact image content uniquely across frequency bands. Our method comprises two key components: a frequency-aware Degradation Estimation Transformer (Dformer) and a Degradation-Adaptive Restoration Transformer (Rformer).
Dformer is proposed to estimate degradation representation, as the degradation prior is not available in the all-in-one image restoration task. Traditional degradation estimation methods [1, 16] often assume a predefined degradation type and estimate degradation level, which makes them less effective in scenarios with multiple unknown degradations. Li et al. [15] suggest obtaining degradation representation using a contrastive learning framework, while Park et al. [17] propose learning a degradation classifier to estimate the type of degradation. Potlapalli et al. [18] utilize prompts to encode degradation-specific information. Unlike these methods, our Dformer captures the essential characteristics of various degradations by decomposing features into different frequency components. By understanding how degradations affect these frequency components, Dformer learns robust priors that effectively guide the restoration process.
Rformer functions as a restoration network. The key challenge in designing such a network lies in developing a dynamic module that adapts to various degradation tasks using guidance from degradation representations. Establishing the correlation between the dynamic module and degradation representation is particularly challenging. Li et al. [15] argue that different degradation tasks necessitate different receptive fields within the restoration network. They designed a dynamic module to adjust the receptive field based on the degradation representation. Park et al. [17] introduced an adaptive discriminative filter-based model to explicitly disentangle the restoration network for multiple degradations. Potlapalli et al. [18] proposed a prompt interaction module to enable dynamic interaction between input features and degradation prompts for guided restoration. In contrast, we recognize that different degradation tasks require the restoration model to focus on distinct frequency components of the degraded image. Rformer adapts to these tasks by employing a degradation-adaptive self-attention mechanism, which allows it to adaptively focus on the most affected frequency components, leading to enhanced restoration performance.
To validate the effectiveness of Dformer and Rformer, we conduct extensive experiments. The results demonstrate that our approach surpasses existing methods across four representative restoration tasks: denoising, deraining, dehazing, and deblurring. Furthermore, our method excels in handling spatially variant degradations and previously unseen degradation levels, highlighting its versatility and robustness.
II Related Works
II-A Multiple degradations image restoration
Numerous restoration methods have been developed for specific tasks, utilizing convolutional neural networks [19, 1, 2, 5, 6, 9, 10] or vision transformers [20, 21, 22, 23, 24, 25, 26]. However, these approaches often struggle to generalize beyond particular types and severities of image degradation. To address this limitation, multi-task and all-in-one methods have been proposed, aiming to handle a wider range of degradation types and levels more effectively.
Multi-task methods [13, 14] focus on training a single model to address multiple image restoration tasks simultaneously by incorporating separate modules for each task in parallel at the input and output layers. For example, Chen et al. [13] developed distinct heads and tails for various tasks, with only the backbone being shared among them. Li et al. [14] introduced a task-specific feature extractor to extract common clean features for different adverse weather conditions. However, these methods still rely on specific degradation priors and are unable to handle unknown degradations.
All-in-one methods [15, 18, 27, 28, 29] aim to address a broad spectrum of image restoration tasks with a single, unified model. Unlike multi-task methods, these approaches eliminate the need for prior knowledge of specific degradations or task-specific designs, making them more versatile and efficient in handling various types of image degradation. Wei et al. [30] and Li et al. [15] pioneered this approach by introducing a new method that utilizes contrastive learning to extract degradation representations, thereby guiding the restoration process. Potlapalli et al. [18] proposed a universal and efficient plugin module that employs adjustable prompts to encode degradation-specific information without prior information on the degradations. Park et al. [27] introduced an adaptive discriminant filter-based degradation classifier to explicitly disentangle the network for multiple degradations.
Unlike the methods discussed earlier, which primarily operate in the spatial domain, this paper presents an all-encompassing image restoration algorithm that carefully considers the variations in frequency across different tasks, aiming to deliver superior results.
II-B Frequency-aware image restoration
Numerous approaches have emerged to address low-level vision problems, with a focus on frequency analysis. Frequency domain frameworks [31, 32, 33, 34, 35, 36] aim to bridge frequency gaps between sharp and degraded images. For instance, Yang et al. [32] use discrete wavelet transforms to facilitate edge feature extraction. Mao et al. [33] distinguish between blurry and sharp images by processing low- and high-frequency components separately using Fast Fourier Transform. Cui et al. [34] propose a selective frequency module that dynamically separates feature maps into distinct frequency components with learnable filters.
Recent studies [37, 38, 39] have explored biases in frequency domain modules. For instance, the self-attention mechanism in transformers acts as a low-pass filter, while CNN convolutions behave like high-pass filters. This underscores the importance of frequency separation to mitigate model biases by handling different frequencies separately. Our study examines varying frequency objectives across image restoration tasks. Denoising and deraining focus on suppressing high-frequency noise, whereas dehazing and deblurring restore high-frequency details. By addressing inherent frequency biases in transformers’ self-attention modules, we propose a frequency-aware all-in-one image restoration method.
III Method
In this section, we present a new all-in-one image restoration method from a frequency perspective, leveraging advanced vision transformers. Our method comprises two main components: a frequency-aware degradation representation learning transformer (Dformer) and a degradation-adaptive Restoration transformer (Rformer). The Dformer captures the essential characteristics of various degradations by decomposing inputs into different frequency components. By understanding how degradations affect these frequency components, Dformer learns robust priors that effectively guide the restoration process. The Rformer employs a degradation-adaptive self-attention module to adaptively focus on the most affected frequency bands, guided by the acquired degradation representations. This adaptive focus is crucial, as different types of degradations impact image content at various frequency bands.
Formally, given an RGB degraded image $\boldsymbol{I}\in\mathbb{R}^{3\times H\times W}$, its degradation representation $d$ can be obtained using $d=\Phi_{D}(\boldsymbol{I})$ where $\Phi_{D}$ denotes the Dformer. Then the retorted image $\hat{\boldsymbol{I}}$ is obtained with $\hat{\boldsymbol{I}}=\Phi_{R}(\boldsymbol{I},d)$ where $\Phi_{R}$ represents the Rformer. In the following sections, we elaborate on the architectures and optimizations of Dformer $\Phi_{D}$ and Rformer $\Phi_{R}$. The overview of the proposed approach is illustrated in Fig. 1 (a).
(a) Frequency-aware dynamic transformers
(b) FA-TB
(c) DA-SA
Figure 1: Overview of the proposed methods. Dformer learns degradation representation and guides Rformer to achieve all-in-one restoration. Input Frequency Decomposition module utilizes DFT and IDFT processes to decompose the input image into multiple frequency-band images. Input Projection module employs a convolution layer to project the input images into the feature maps. Frequency-Aware Transformer Blocks (FA-TB) is detailed in (b). Output Projection module includes 2D average pooling and two-layer MLP to refine and project degradation representation. Degradation Projection includes a two-layer MLP. The architecture of Rformer follows Uformer, but employs a new degradation-adaptive self-attention mechanism (DA-SA), as detailed in (c), to adaptively handle varying levels and types of image degradation.
III-A Frequency-aware degradation representation learning transformer
We propose Dformer, a frequency-aware transformer specifically designed to learn degradation representations by accounting for the differences in how various degradation types affect image content across frequency bands. Dformer constructs a hierarchical encoder network following the architecture of the Swin Transformer [40], as illustrated in Fig. 1 (a). Dformer incorporates two key designs: 1) Input Frequency Decomposition Module decomposes the input degraded image into distinct frequency bands. 2) Frequency-aware Swin Transformer Block performs self-attention both within and between these frequency bands, effectively learning degradation representations.
Input frequency decomposition module. Given the RGB degraded image $\boldsymbol{I}\in\mathbb{R}^{3\times H\times W}$, the module first performs a 2D discrete Fourier transform (DFT) to obtain the Fourier spectrum of $\boldsymbol{I}$. Then the Fourier spectrum of $k$-th frequency band, denoted as $\operatorname{F-Band}_{k}(\boldsymbol{I})\in\mathbb{C}^{H\times W}$, can be obtained by:
$\begin{split}\operatorname{F-Band}_{k}(\boldsymbol{I})=\begin{cases}{{\mathcal{F}(\boldsymbol{I})_{ij},}}&{{\mathrm{if~{}}|i-\lfloor{\frac{n}{2}}\rfloor|,|j-\lfloor{\frac{n}{2}}\rfloor|\in[l_{k},r_{k}]}}\\ {{0,}}&\mathrm{otherwise}\\ \end{cases},\end{split}$ (1)
where $\mathcal{F}:\mathbb{R}^{n\times n}\to\mathbb{C}^{n\times n}$ denote the 2D DFT. $l_{k}$ and $r_{k}$ denote the minimum and maximum frequencies for each band, respectively. The frequency range is divided into $L$ bands, where the first band only contains the direct current (DC) component (i.e., $l_{1}=r_{1}=0$), and the remaining bands divide the entire frequency range equally. The Fourier spectrum of each frequency band is transformed back to the spatial domain using the 2D inverse DFT, denoted by ${\mathcal{F}}^{-1}:\mathbb{C}^{n\times n}\to\mathbb{R}^{n\times n}$:
$\begin{split}\boldsymbol{I_{k}}={{\mathcal{F}}^{-1}(\operatorname{F-Band}_{k}(\boldsymbol{I})).}\end{split}$ (2)
where ${\mathcal{F}}^{-1}:\mathbb{C}^{n\times n}\to\mathbb{R}^{n\times n}$ denote the 2D inverse DFT.
This module generates $L$ new images $\{I_{1},I_{2},\ldots,I_{L}\}$, each corresponding to a different frequency band of the degraded image $I$. These images are then passed through a shared $3\times 3$ convolutional layer to extract low-level features. The extracted features are subsequently processed through $K$ shared encoder stages. Each stage consists of $N$ frequency-aware Swin Transformer blocks and a downsampling layer, except for the last stage. After the $K$ encoder stages, we use an output projection, which includes 2D average pooling and two-layer MLP, to generate a degradation representation vector. Next, we will detail the design of our frequency-aware Swin Transformer blocks, which are specifically tailored to capture degradation representations for fully considering every frequency band.
Frequency-aware Transformer block. The widely used Swin Transformer block employs a shifted window-based self-attention mechanism to efficiently capture both local and global contextual information. Unlike the original Swin Transformer block, which processes a single input image $I$, our enhanced block processes $L$ input images, $\{I_{1},I_{2},\ldots,I_{L}\}$, derived from the input frequency decomposition module. To enable the Swin Transformer block to handle multiple frequency-band inputs and fully leverage their contents, we introduce a new frequency-aware transformer block, as illustrated in Fig. 1 (b). This block incorporates new designs in self-attention mechanisms, positional encoding, and masking techniques.
We introduce Intra- and Inter-Band shifted window-based self-attention mechanisms to facilitate adaptive interactions within and between frequency bands. Intra-band self-attention facilitates interactions among distinct pixels within each frequency band, essentially performing self-attention computations independently for each band within the Swin Transformer block. This method ensures complete isolation between different frequency bands, focusing exclusively on intra-band interactions. On the other hand, inter-band self-attention explicitly manages interactions across different frequency bands. Utilizing a window-based strategy, it computes self-attention between pixels from different frequency bands within the same spatial window. This approach allows for a more detailed examination of frequency disparities within localized regions.
To adapt the relative positional encoding and window shifting mechanism within the Swin Transformer block to variations in token count and dimensions, we propose integrating a one-dimensional absolute frequency domain positional encoding alongside the original two-dimensional relative spatial positional encoding. Additionally, to facilitate the window shifting mechanism, we introduce an enhanced masking mechanism. This ensures that interactions occur exclusively among tokens within spatially adjacent shifted windows that meet the frequency criteria for both intra- and inter-band self-attention.
III-B Degradation-adaptive restoration transformer
After obtaining the degradation representations, we incorporate them into a restoration transformer (Rformer), as illustrated in Fig. 1 (a). The architecture of Rformer follows Uformer [23], but employs a new degradation-adaptive self-attention mechanism to adaptively focus on the most affected frequency bands, guided by the acquired degradation representations. We illustrate the degradation-adaptive self-attention mechanism in Fig. 1 (c).
Let $\boldsymbol{z}$ be the self-attention map in the transformer block. The frequency band of $\boldsymbol{z}$ can be obtained by Eq. 1. $\operatorname{F-Band}_{k}(\boldsymbol{z})$ denotes the $k$-th frequency band partitioned from the attention maps. After frequency decomposition, the frequency scaling is performed as follows:
$\begin{split}\boldsymbol{z}^{{}^{\prime}}=\boldsymbol{z}+\sum_{k>1}^{L}\boldsymbol{M}_{k-1}{\mathcal{F}}^{-1}(\operatorname{F-Band}_{k}(\boldsymbol{z})),\end{split}$ (3)
where $\boldsymbol{z}^{{}^{\prime}}$ represent the rescaled attention map, and $\boldsymbol{M}_{k}$ denotes the scaling coefficient for the $k$-th frequency band. The direct component of $z$ serves as a baseline, remaining unscaled to provide a reference for scaling other frequency bands. The set of scaling coefficients, $\boldsymbol{M}=\{\boldsymbol{M}_{1},\ldots,\boldsymbol{M}_{L-1}\}$, is learned through a degradation projection process implemented by a two-layer MLP. This MLP takes the degradation representations $d$, derived from Dformer, as input. The MLP is initialized such that the values of $M$ are zero, resulting in $\boldsymbol{z}^{\prime}=\boldsymbol{z}$. Essentially, the attention map $z$ is decomposed into multiple frequency bands, each of which is scaled by a coefficient learned through a degradation-aware projection. This allows for adaptive restoration based on the degradation characteristics.
III-C Composite training loss
The training of our approach is carried out in two distinct stages. Initially, Dformer is trained to learn degradation representations with a contrastive learning loss $\mathcal{L}_{cl}$. We consider $\boldsymbol{d}$ as the degradation representation of the anchor sample, $\boldsymbol{d}^{+}$ and $\boldsymbol{d}^{-}$ as the degradation representation of positive and negative samples obtained through the MoCo framework, where positive samples and the anchor sample come from the same degraded image, while negative samples come from other degraded images. $\mathcal{L}_{cl}$ is defined by:
$\begin{split}\mathcal{L}_{cl}=-\operatorname{log}\frac{\operatorname{exp}(\boldsymbol{d}\cdot\boldsymbol{d}^{+}/\tau)}{\sum_{\boldsymbol{d}^{-}\in Queue}\operatorname{exp}(\boldsymbol{d}\cdot\boldsymbol{d}^{-}/\tau)}\end{split}$ (4)
where $Queue$ represents the negative sample queue in the MoCo framework, and $\tau$ denotes the temperature hyperparameter.
In the second stage, we train the Dformer and Rformer together by using a composite loss function. This loss function comprises two distinct components:
$\mathcal{L}=\mathcal{L}_{cl}+\mathcal{L}_{rec},$ (5)
where $\mathcal{L}_{rec}=\frac{1}{T}\sum_{i=1}^{T}|\hat{I}_{i})-y_{i}|$ is a L1 loss. Here, $\hat{I}_{i}$ denotes the recovered image through Rformer, and $y$ is the corresponding clean image.
IV Experiments and Results
In this section, we comprehensively evaluate and analyze our methods across four tasks: denoising, deraining, dehazing, and deblurring.
IV-A Experimental Setup
Datasets. Following the existing works [15, 18], we assess the effectiveness of the proposed approaches in multi-degradation restoration using seven datasets: BSD400 [41], BSD68 [41], WED [42], and Urban100 [43] for image denoising, Rain100L [44] for image deraining, RESIDE [45] for image dehazing, and GoPro [46] for image deblurring.
Implementation details. The training settings are outlined following AirNet [15]. AdamW is used as the optimizer. The training phase consists of $1000$ epochs: the first $100$ epochs train the Encoder with Contrastive Loss optimization for warm-up, and the remaining $900$ epochs optimize the entire network using total loss optimization. The learning rate starts at $43e-$ 4, reducing to $53e-$ 5 after $60$ epochs, and then starts at $41e-$ 4 for the remaining epochs, halving every $125$ epochs. We fix the image patch size at $128\times 128$ and apply random data augmentations. The batch size is set to $400\times N$, where $N$ is the number of degradation types. We set the number of frequency bands to $L=2$ to balance efficiency and performance, as demonstrated in Section VII.
Metrics. In line with Li et al. [15], we employ two widely used metrics for quantitative comparisons: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity (SSIM). A superior performance is indicated by higher values of these metrics.
IV-B Comparison to the state-of-the-art
We first perform a comparison to the state-of-the-art in the conventional ”noise-rain-haze” setting to showcase the superiority of our approach. Our comparison encompasses four single-degradation image restoration techniques, namely BDRNet [47], LP-Net [48], FDGAN [49], and MPRNet [50], alongside the multi-task method for multiple degradation image restoration, DL [51]. We are also specifically evaluating two specialized all-in-one methods, AirNet [15] and PromptIR [18].
The results, as shown in Table I, highlight the superiority of all all-in-one methods over other baselines for single degradation, underscoring their ability to address various unknown degradations within a unified framework. Notably, our approach demonstrates even better performance compared to other all-in-one methods. Specifically, we surpass AirNet [15] across all tasks, achieving an average performance improvement of $1.12$ dB PSNR and $0.011$ SSIM. Furthermore, we outperform PromptIR [18] in denoising and deraining tasks, with an average performance improvement of $0.26$ dB PSNR and $0.008$ SSIM.
Qualitative examples are presented in Fig. 2. Compared to AirNet [15] and PromptIR [18], our approach better preserves edge details when performing denoising and deraining, and achieves better color fidelity during dehazing.
TABLE I: Comparison to the state-of-the-art on the conventional “noise-rain-haze” setting. Existing all-in-one methods surpass other baselines designed for single degradation tasks, whereas our approach achieves superior performance.
Method Denoise Derain Dehaze Average
BSD68 $(\sigma=15)$ BSD68 $(\sigma=25)$ BSD68 $(\sigma=50)$ Rain100L SOTS
BRDNet[47] 32.26/0.898 29.76/0.836 26.34/0.693 27.42/0.895 23.23/0.895 27.80/0.843
LPNet[52] 26.47/0.778 24.77/0.748 21.26/0.552 24.88/0.784 20.84/0.828 23.64/0.738
FDGAN[49] 30.25/0.910 28.81/0.868 26.43/0.776 29.89/0.933 24.71/0.929 28.02/0.883
MPRNet[50] 33.54/0.927 30.89/0.880 27.56/0.779 33.57/0.954 25.28/0.955 30.17/0.899
DL[51] 33.05/0.914 30.41/0.861 26.90/0.740 32.62/0.931 26.92/0.931 29.98/0.876
AirNet[15] 33.92/0.933 31.26/0.888 28.00/0.797 34.90/0.968 27.94/0.962 31.20/0.910
PromptIR[18] 33.98/0.933 31.31/0.888 28.06/0.799 36.37/0.972 30.58/0.974 32.06/0.913
Ours 34.59/0.941 31.83/0.900 28.46/0.814 37.50/0.980 29.20/0.972 32.32/0.921
(a) Degraded images
(b) AirNet
(c) PromptIR
(d) Ours
(e) Groundtruth
Figure 2: The performance of various methods on denoising of $\sigma=25$ (first row), deraining (second row), and dehazing tasks (last row). In the blue-highlighted regions, our method demonstrates superior edge detail preservation for both deraining and denoising tasks. Additionally, it achieves better color fidelity in the dehazing task compared to other methods.
IV-C Comparison on the number of degradation types
In this experiment, we conduct a comparative analysis between the proposed method, AirNet [15], and PromptIR [18] across different numbers of degradations to assess the stability of our approach. The experimental results in Table II show that as the number of degradation types increases, the network’s ability to restore images diminishes, resulting in a performance decline. Notably, both AirNet and PromptIR experience clear performance degradation when tasked with handling multiple degradations simultaneously. For instance, the PSNR for deraining drops from 38.31 dB to 34.7 dB for AirNet, and from 39.32 dB to 36.14 dB for PromptIR, as the number of combined degradation types increases from 2 to 4. This decline in performance occurs due to potential conflicts between different tasks during joint learning, which AirNet and PromptIR struggle to manage effectively. In contrast, our method explicitly addresses task disparities in frequency domain through frequency-aware dynamic transformers. Consequently, our method experiences a milder drop of only 1.43 dB, from 39.51 dB PSNR to 37.35 dB PSNR, showcasing superior stability across varying numbers of degradation types.
TABLE II: Comparison on the number of degradation types. As the number of combined degradation types increases, our proposed approach demonstrates superior performance stability compared to AirNet and PromptIR.
D-Types Method Denoise Derain Dehaze Deblur
BSD68 $(\sigma=15)$ BSD68 $(\sigma=25)$ BSD68 $(\sigma=50)$ Rain100L SOTS GOPro
1 AirNet 34.14/0.936 31.49/0.893 28.23/0.806 - - -
PromptIR 34.34/0.940 31.71/0.900 28.49/0.813 - - -
Ours 34.74/0.943 31.98/0.903 28.66/0.820 - - -
2 AirNet 34.11/0.935 31.46/0.892 28.19/0.804 38.31/0.982 - -
PromptIR 34.26/0.937 31.61/0.895 28.37/0.810 39.32/0.986 - -
Ours 34.69/0.942 31.93/0.902 28.59/0.818 39.51/0.989 - -
3 AirNet 33.92/0.933 31.26/0.888 28.01/0.798 34.90/0.968 27.94/0.961 -
PromptIR 33.98/0.933 31.31/0.888 28.06/0.799 36.37/0.972 30.58/0.974 -
Ours 34.59/0.941 31.83/0.900 28.46/0.814 37.50/0.980 29.20/0.972 -
4 AirNet 33.89/0.932 31.21/0.887 27.97/0.795 34.70/0.964 27.41/0.956 26.36/0.799
PromptIR 33.91/0.933 31.24/0.888 28.01/0.797 36.14/0.968 29.82/0.969 27.16/0.820
Ours 34.58/0.941 31.83/0.900 28.46/0.813 37.35/0.980 28.93/0.971 27.42/0.829
IV-D Results on various combined degradations
In this section, we examine the impact of various combinations of degradation types on model performance, as detailed in Table III. When we randomly select and combine two out of the four degradation types, we observe that denoising performance remains relatively stable, regardless of whether it is combined with deraining, dehazing, or deblurring. This stability is likely because the denoising task dominates the training process, benefiting from a larger dataset across three noise levels: $\sigma=15$, $\sigma=25$, and $\sigma=50$.
For the deraining task, performance is optimal when combined with denoising, compared to combinations with dehazing or deblurring. This is likely due to both deraining and denoising focusing on recovering high-frequency details, thus aligning their frequency optimization directions. Conversely, dehazing and deblurring aim to remove low-frequency content, and their performance is enhanced when combined with denoising, due to the larger training dataset. Similar trends are observed when we randomly select and combine three out of the four degradation types, further supporting these findings.
TABLE III: Results on various combined degradations. Tasks can enhance each other when their degradation types (e.g., deraining and denoising) share similar frequency optimization directions. Conversely, when degradation tasks (e.g., deraining and dehazing) have conflicting optimization goals, a performance drop is observed.
Degradation Denoise Derain Dehaze Deblur
Noise Rain Haze blur BSD68 $(\sigma=15)$ BSD68 $(\sigma=25)$ BSD68 $(\sigma=50)$ Rain100L SOTS GOPro
$\checkmark$ $\checkmark$ 34.69/0.942 31.93/0.902 28.59/0.818 38.93/0.984 - -
$\checkmark$ $\checkmark$ 34.66/0.942 31.91/0.902 28.56/0.818 - 29.01/0.972 -
$\checkmark$ $\checkmark$ 34.67/0.942 31.92/0.902 28.57/0.817 - - 29.05/0.871
$\checkmark$ $\checkmark$ - - - 36.55/0.976 28.64/0.971 -
$\checkmark$ $\checkmark$ - - - 37.99/0.981 - 28.69/0.863
$\checkmark$ $\checkmark$ - - - - 28.02/0.968 26.74/0.809
$\checkmark$ $\checkmark$ $\checkmark$ 34.59/0.941 31.83/0.900 28.46/0.814 37.50/0.980 29.20/0.972 -
$\checkmark$ $\checkmark$ $\checkmark$ 34.65/0.942 31.89/0.901 28.54/0.816 38.72/0.984 - 28.99/0.870
$\checkmark$ $\checkmark$ $\checkmark$ 34.62/0.942 31.87/0.901 28.52/0.815 - 28.65/0.970 28.23/0.851
$\checkmark$ $\checkmark$ $\checkmark$ - - - 36.03/0.974 28.15/0.968 26.70/0.809
IV-E Ablation Studies
In this section, we present the ablation experiments outlined in Table IV-E and IV-E to validate the effectiveness of the proposed Dformer and Rformer models, along with their individual components. These experiments are conducted under the standard ”noise-rain-haze” setting. For clarity and conciseness, we report only the average PSNR and SSIM metrics.
Dformer. The key components of Dformer are the Input Frequency Decomposition (IFD) and Frequency-Aware Transformer Blocks (FA-TB). For comparison, we used the Swinformer as a baseline, which incorporates standard Swin Transformer blocks. Additionally, we created a second baseline by combining our Input Frequency Decomposition (IFD) with Swinformer. Both baselines, along with our Dformer, utilize the Rformer for restoration. As shown in Table IV-E, incorporating IFD into the Swinformer results in a slight performance improvement, with the average PSNR increasing from 31.53 to 31.63. However, when replacing the standard Swin Transformer blocks with our FA-TB, the average PSNR further improves from 31.63 to 32.32. These results underscore the importance of IFD and FA-TB in learning better degradation representations and enhancing overall image restoration performance.
Rformer. Rformer is designed following the architecture of Uformer, so we use Uformer as the primary baseline. Uformer addresses all tasks simultaneously without leveraging any degradation priors. The key component of Rformer is the Degradation-Adaptive Self-Attention (DA-SA). DA-SA dynamically rescales different frequency bands of the attention map to achieve adaptive restoration, guided by degradation representations acquired from Dformer. To further evaluate DA-SA, we developed an additional baseline where the rescaling is performed using learnable parameters without any degradation guidance. As shown in Table IV-E, our Rformer achieves the best performance, demonstrating the importance of DA-SA in enhancing restoration by effectively adapting to degradation characteristics.
TABLE IV: Ablation study
of Dformer.
Methods Average Swinformer 31.53/0.914 Swinformer+IFD 31.62/0.915 Dformer 32.32/0.921
TABLE V: Ablation study
of Rformer.
Methods Average Uformer 31.01/0.902 Scaling Uformer 30.91/0.898 Rformer 32.32/0.921
IV-F Further analysis
Performance on spatially variant degradation. We analyze the performance of the proposed method under spatially variant degradation, aiming to highlight its enhanced capability in restoring spatial heterogeneity. Following the experimental setup of AirNet [15], we partition each clean image of the BSD68 [41] dataset into four regions. Subsequently, Gaussian noise with $\sigma\in\{0,15,25,50\}$ is injected into each region individually to create a new test set. We then assess the model trained solely on the standard denoising task using this new test set. As shown in Table VI, our method outperforms both AirNet [15] and PromptIR [18], achieving a PSNR improvements of 0.34 dB over AirNet and 0.11 dB over PromptIR.
TABLE VI: Performance on spatially variant degradation. Under spatially variant degradation, our method showcases superior denoising performance compared to existing methods.
Method Denoise
BSD68 $(\sigma\in\{0,15,25,50\})$
AirNet 31.42/0.892
PromptIR 31.65/0.899
Ours 31.76/0.902
TABLE VII: Generalization to unseen degradation level. Our method achieves superior generalization performance over the existing AirNet and PromptIR.
Method Denoise
BSD68 $(\sigma\in[15,25])$ BSD68 $(\sigma\in[25,50])$
AirNet 31.80/0.887 28.30/0.782
PromptIR 32.34/0.908 29.18/0.830
Ours 33.13/0.918 29.34/0.832
Generalization to unseen degradation levels. To analyze the generalization capability of our model for unseen degradation levels, we evaluate it on the BSD68 [41] test set. Specifically, our model, trained solely on $\sigma\in\{15,25,50\}$, is tested with randomly sampled values from the ranges $\sigma\in[15,25]$ and $\sigma\in[25,50]$. The results shown in Table VII highlight the superior generalization performance of our model over AirNet and PromptIR.
The effect of the number of frequency bands. Finally, we examine the impact of the number of frequency bands, denoted as $L$, on both efficiency and performance. Specifically, we compare the performance between $L=2$ and $L=3$ under the standard ”noise-rain-haze” setting, as detailed in Table VIII. The results show that increasing the number of frequency bands from $L=2$ to $L=3$ improves restoration performance across all three tasks. This aligns with our expectations, as a higher value of $L$ enables the model to more finely distinguish between different degradation types at various frequencies, enhancing overall performance. However, the number of tokens in Intra- and Inter-Band attention scales with $L$, leading to attention maps and time complexity increasing proportionally. For example, the training time per epoch for Dformer is 70 seconds for $L=2$ and 90 seconds for $L=3$. To balance efficiency and performance, we have chosen to use $L=2$ as the default value for all experiments.
TABLE VIII: The effect of frequency decomposition on three degradation types. The restoration performance for all three tasks boosts when the number of frequency bands is increased from $L=2$ to $L=3$, while efficiency decreases
Method Denoise Derain Dehaze Training time of Dformer
BSD68 $(\sigma=15)$ BSD68 $(\sigma=25)$ BSD68 $(\sigma=50)$ Rain100L SOTS
L=2 34.59/0.941 31.83/0.900 28.46/0.814 37.50/0.980 29.20/0.972 70s/epoch
L=3 34.61/0.944 31.92/0.902 28.54/0.816 37.88/0.982 29.33/0.974 90s/epoch
V Conclusion
This work presents an all-in-one image restoration model leveraging advanced vision transformers, inspired by the fact that various degradations uniquely impact image content across different frequency bands. The model consists of two primary components: the frequency-aware Degradation Prior Learning Transformer (Dformer) and the Degradation-Adaptive Restoration Transformer (Rformer). The Dformer captures degradation representations by using an input frequency decomposition module and frequency-aware Swin Transformer blocks. Guided by these learned representations, the Rformer utilizes a degradation-adaptive self-attention module to selectively focus on the most affected frequency components for restoration. Extensive experimental results demonstrate the superiority of our approach over existing methods in four key restoration tasks: denoising, deraining, dehazing, and deblurring. Furthermore, our method excels in handling spatially variant degradations and previously unseen degradation levels. These findings underscore the potential of our frequency-based perspective and advanced transformer design to significantly advance the field of image restoration.